1. How much 84 octane gas and 98 octane gas should be blended to make 14 gallons of 92 octane gas?
2. A train leaves a station and travels north at 50 km/h. Two hours later a second train leaves on a parallel track traveling north at 90 km/h. how far from the station will they meet?
3.

Answer by checkley77(12844)

98x+84(14-x)=92*14
98x+1176-84x=1288
14x=1288-1176
14x=112
x=112/14
x=8 gallons of 98 octane gas is used.
14-8=6 gallons of 84 octane gas is used.
98*8+84*6=92*14
784+504=1288
1288=1288
-------------------------------------------------
50t=90(t-2)
50t=90t-180
50t-90t=-180
-40t=-180
t=-180/-40
t=4.5 hours after the first train leaves the station they will meet.
50*4.5=225 miles is the distance traveled.
90(4.5-2)=90*2.5=225 miles.